# On homology spheres of the trivial local equivalence class

### v1

@@ Abstract.
In the homology cobordism group $\Theta_{\mathbb{Z}}^{3}$, it is not known if there is a non-trivial linear dependence between Seifert fibered spheres. Based on involutive Heegaard Floer theory, Hendricks, Manolescu, and Zemke introduced the local equivalence group $\mathfrak{I}$ with the homomorphism $h:\Theta_{\mathbb{Z}}^{3}\rightarrow\mathfrak{I}$. And it is possible to find a non-trivial linear dependence between the image of Seifert fibered spheres in $\mathfrak{I}$ using the work of Dai and Stoffregen. Therefore, it is interesting to ask if such a linear dependence in $\mathfrak{I}$ originates from $\Theta_{\mathbb{Z}}^{3}$. In this paper, we prove that this is not the case in general. We show that infinitely many homology spheres linearly dependent in $\mathfrak{I}$ are linearly independent even in the rational homology cobordism group $\Theta_{\mathbb{Q}}^{3}$, by employing the $r_{s}$-invariants from the filtered instanton Floer homology developed by Nozaki, Sato, and Taniguchi. Moreover, these homology spheres include homology spheres of the trivial local equivalence class of $\operatorname{Pin(2)}$-equivariant Seiberg–Witten Floer stable homotopy type.
@@ 1. Introduction
The homology cobordism group $\Theta^{3}_{\mathbb{Z}}$ of homology spheres, introduced by González-Acuña [GA70], has played a central role in the development of low-dimensional topology over the last five decades [Man18, Hom23, Şav24]. The algebraic structure of $\Theta^{3}_{\mathbb{Z}}$ is linked to several topological problems such as the three fibers conjecture [FS87, Law88], the disproof of the triangulation conjecture for topological manifolds in high dimensions [Man16], and so on. While it is known that the homology cobordism group contains an infinite rank summand by [DHST23], the existence of a torsion element is one of the most notable open problems about the structure of $\Theta^{3}_{\mathbb{Z}}$.
Seifert fibered spheres have been core objects to study the structure of $\Theta^{3}_{\mathbb{Z}}$. In particular, the subgroup $\Theta^{3}_{SF}$ generated by Seifert fibered spheres itself contains an infinite rank subgroup in $\Theta^{3}_{\mathbb{Z}}$ [Fur90, FS90]. Moreover, the aforementioned infinite rank summand [DHST23] lies also in $\Theta_{SF}^{3}$. On the other hand, there are several infinite families of Seifert fibered spheres known to bound contractible manifolds or homology balls [AK79, CH81, FS81, Fic84, KŞ25], namely trivial in $\Theta_{\mathbb{Z}}^{3}$. However, it is neither known in $\Theta_{\mathbb{Z}}^{3}$ whether there exists a non-trivial Seifert fibered sphere of finite order, nor, more generally, whether there exists any non-trivial linear dependence among them. More precisely, we ask:
@@
Let $\mathfrak{I}_{SF}$ be the image of $\Theta_{SF}^{3}$ under $h$. While $\mathfrak{I}_{SF}$ has been particularly studied in [DM19, DS19] and is known to be isomorphic to $\mathbb{Z}^{\infty}$, it is possible to find many linear dependences in $\mathfrak{I}_{SF}$. Therefore, this fact and Question 1.1 motivate to study the kernel of the restriction of $h$ to $\Theta^{3}_{SF}$. In this paper, we detect a substantial structural difference between $\Theta^{3}_{SF}$ and $\mathfrak{I}_{SF}$, providing affirmative evidence for Question 1.1.
Theorem 1.2.
We remark that we obtain linear independence not only in $\Theta_{\mathbb{Z}}^{3}$, but also in the rational homology cobordism group $\Theta_{\mathbb{Q}}^{3}$ of rational homology spheres. Together with the argument in Lemma 3.3, Furuta’s result [Fur90] is sufficient to prove linear independence in $\Theta^{3}_{\mathbb{Z}}$ of our homology spheres, which will be constructed later. However, this does not directly imply linear independence in $\Theta_{\mathbb{Q}}^{3}$, since the kernel of $\Theta_{\mathbb{Z}}^{3}\rightarrow\Theta_{\mathbb{Q}}^{3}$ is known to be non-trivial [FS84, AL18, Şav20, Sim21].
@@
Therefore, potential candidates for a negative answer to Question 1.1 arise from linear dependence relations with respect to this explicit basis $\{h(B(n))\}_{n\geq 0}$ of $\mathfrak{I}_{SF}$. We construct homology spheres in Theorem 1.2 by relying on the aforementioned results [DM19, DS19] in terms of monotone graded subroots.
We consider the families of the Brieskorn spheres:
@@
Their monotone graded subroots and hence their local equivalence classes were computed by Karakurt and the second author [KŞ22]. By using the fact that every monotone graded subroot is decomposed into simpler ones [DS19, Theorem 1.1], we first find each linear expression of the local equivalence class of $Y_{1}(n)$ and $Y_{2}(n)$ with respect to $h(B(n))$. From these linear relations, we obtain infinitely many homology spheres lying in $\operatorname{ker}(h)$ as follows:
$\big{\{}Y_{1}(n)\ \#\ -B(2n)\ \#\ B(n)\ \#\ -nB(0)\big{\}}_{n\geq 1}\cup\big{\{}Y_{2}(n)\ \#\ -B(2n-1)\ \#\ B(n)\ \#\ -nB(0)\big{\}}_{n\geq 1}.$
Note that Theorem 1.2 indicates that the $d$-invariant [OS03a], the $\underline{d\mkern-2.0mu}\mkern 2.0mu$- and $\bar{d}$-invariants [HM17], and the $\phi_{n}$-invariants for all $n\geq 1$ [DHST23] must vanish for homology spheres above. In other words, letting $Y$ be any homology sphere generated by the above set, we see that:
$\underline{d}(Y)=d(Y)=\overline{d}(Y)=\phi_{n}(Y)=0.$
We observe that the $d$-invariant is not a complete rational homology cobordism invariant. Other homology spheres of infinite order in $\Theta^{3}_{\mathbb{Q}}$ with vanishing $d$-invariants can be easily deduced from [NST24]. We also note that the $\bar{\mu}$-invariant [Neu80, Sie80] vanishes by the result of [DS19].
$\Theta^{3}_{\mathbb{Z}}\to\mathfrak{LE},$
where the latter group denotes the local equivalence group of the $\operatorname{Pin}(2)$-equivariant Seiberg–Witten Floer stable homotopy type. Moreover, Dai, Sasahira, and Stoffregen [DSS23] recently proved that two Seifert fibered spheres have the same image in $\mathfrak{LE}$ if and only if they have the same image in $\mathfrak{I}$.
Motivated from this result, we provide another infinite family using the following family of the Brieskorn spheres:
@@
Since its monotone graded subroot is known to be the same as the one of $B(4n)$ by [KŞ22], the resulting homology spheres are as follows:
$\{Y_{3}(n)\ \#\ -B(4n)\}_{n\geq 1}.$
These homology spheres also fall under Theorem 1.2. Moreover, they have the trivial local equivalence class in $\mathfrak{LE}$ as well. In particular, the $\kappa$-invariant [Man14] and the $\kappa o_{i}$-invariants for $i=0,\ldots,7$ [Lin15] from Seiberg–Witten Floer stable homotopy type, and the $\delta$-invariant [Frø10], the $\alpha$-, $\beta$-, and $\gamma$-invariants [Man16], and the $\underline{\delta}$- and $\overline{\delta}$-invariants [Sto17b] from Seiberg–Witten Floer homology must vanish for these homology spheres. Namely, letting $Y$ be any homology sphere generated by the above set, we also have the following:
$\alpha(Y)=\beta(Y)=\gamma(Y)=\underline{\delta}(Y)=\delta(Y)=\overline{\delta}(Y)=\kappa(Y)=\kappa o_{i}(Y)=0.$
To the authors’ best knowledge, it is not known whether there exists a Seifert fibered sphere for Theorem  1.2. The reason is that it is difficult to compute the $r_{0}$-invariant in general when the Fintushel–Stern $R$-invariant [FS85] is negative. Moreover, Issa and McCoy [IM20] proved that for a Seifert fibered sphere $Y$ with $d(Y)=0$, $R(Y)$ is always negative. Therefore, we conclude the introduction with the following question:
Question 1.3.
@@
[…]
@@ Acknowledgments
The authors would like to thank Marco Golla, JungHwan Park, and Masaki Taniguchi for helpful conversations. The authors are also grateful to Hayato Imori and Imogen Montague for reading the draft and providing detailed feedback. The first author is partially supported by the Samsung Science and Technology Foundation (SSTF-BA2102-02) and the NRF grant RS-2025-00542968.
@@ 2. Preliminaries on Floer Theories
The proof of Theorem 1.2 involves two major theories: involutive Heegaard Floer homology [HM17] and filtered instanton Floer homology [NST24]. In this section, we briefly recall the basic concepts, key properties, and essential notions of each theory. For the first two subsections, we review the involutive Heegaard Floer homology and its local equivalence class in terms of monotone graded roots. In the third, we consider analogue in the Seiberg–Witten Floer theory. In the last, we briefly recall the filtered instanton homology and the construction of the $r_{s}$-invariants.
@@ 2.1. Involutive Heegaard Floer homology and local equivalence group
$h:\Theta_{\mathbb{Z}}^{3}\to\mathfrak{I},\ \ h(Y)=(CF^{-}(Y),\iota)[-2],$
@@ 2.2. Monotone graded subroots for Seifert fibered spheres
In this subsection, we explain the roles of graded roots and their monotone graded subroots to describe local equivalence classes of Seifert fibered spheres. We follow the references [Ném05, DM19, DS19, KŞ22].
In [Ném05], Némethi introduced two combinatorial objects to compute Heegaard Floer homology of Seifert fibered spheres effectively. The first object is called a graded root $(R,\nu)$ equipped with a grading function $\nu$ on the vertex set of $R$, where $R$ is an upwards-opening tree with an infinite downwards stem defined from the negative definite plumbing graph for a Seifert fibered sphere. The second one is a graded $\mathbb{F}[U]$-module, called the lattice homology $\mathbb{H}^{-}(R)$. It is combinatorially defined from $(R,\nu)$, whose grading is induced by $\nu$ with the formal variable $U$ of grading $2$. The grading on $\mathbb{H}^{-}(R)$ is denoted by the same symbol $\nu$ by abuse of notation. See [Ném05, DM19] for precise definitions.
@@
By combining the results of Ozsváth and Szabó [OS03b], Némethi [Ném05], and Dai and Manolescu [DM19], we have the following graded $\mathbb{F}[U]$-module isomorphism:
@@
where $\sigma$ denotes a grading shift [^1] For a Seifert fibered sphere, this shift is determined by algebraic topology of the minimal negative definite plumbing bounded by $Y$. See [KŞ22, Section 2.3] and references therein for a precise definition. and $\iota_{*}$ is the induced involution by $\iota$ on $HF^{-}(Y)$.
Following [DM19], we next construct a free and finitely generated $\mathbb{F}[U]$-chain complex whose homology is the same as $\mathbb{H}^{-}(R)$. See [DM19, Lemma 4.3] for a proof of this fact.
@@
Then the generators of $C_{*}(R)$ are given as follows. For each leaf $v_{i}$, we place a single generator (denoted by also $v_{i}$ by abuse of notation) in grading $\nu(v_{i})$, so that we introduce an entire tower of generators $\mathbb{F}[U]v_{i}$. For each angle $\alpha_{i}$, we similarly place a single generator (denoted by also $\alpha_{i}$ by abuse of notation) in grading $\nu(\alpha_{i})+1$. We next define our differential to be identically zero on $v_{i}$, and set
@@
Let $(R,\nu,J)$ be a symmetric graded root, and let $\mathcal{G}_{R}$ be the image of the grading function $\nu:R\to\mathbb{Z}$. For a positive integer $n$, let $h_{1},\ldots,h_{n}$ (resp. $r_{1},\ldots,r_{n}$) be a sequence of decreasing (resp. increasing) even integers in $\mathcal{G}_{R}$ such that $h_{n}\geq r_{n}$. Then a monotone graded subroot
@@
By construction, the standard complex $(C_{*}(M),J)$ of a monotone graded subroot $M$ is a subcomplex of $(C_{*}(R),J)$, and it also forms an $\iota$-complex itself, see [DM19, Section 6]. However, it is sufficient to describe the local equivalence class of a Seifert fibered sphere completely, as proven by Dai and Manolescu.
@@
Now we present a key theorem, provided by Dai and Stoffregen [DS19], to prove Theorem 1.2. It describes how a monotone graded subroot decomposes into simpler ones up to local equivalence.

### v2

@@ Abstract.
In the homology cobordism group $\Theta_{\mathbb{Z}}^{3}$, it is not known if there are non-trivial linear dependences between Seifert fibered spheres. Based on involutive Heegaard Floer theory, Hendricks, Manolescu, and Zemke introduced the local equivalence group $\mathfrak{I}$ along with the homomorphism $h:\Theta_{\mathbb{Z}}^{3}\rightarrow\mathfrak{I}$. Using the work of Dai and Stoffregen, one can find non-trivial linear dependences between the images of Seifert fibered spheres under $h$. Therefore, it is interesting to ask if such dependences in $\mathfrak{I}$ originate from $\Theta_{\mathbb{Z}}^{3}$. In this paper, by employing the $r_{s}$-invariants from the filtered instanton Floer homology developed by Nozaki, Sato, and Taniguchi, we provide certain conditions to guarantee that such relations are not realized even in the rational homology cobordism group. We also discuss the local equivalence class of the $\operatorname{Pin(2)}$-equivariant Seiberg–Witten Floer stable homotopy type.
@@ 1. Introduction
The homology cobordism group $\Theta^{3}_{\mathbb{Z}}$ of homology spheres, introduced by González-Acuña [GA70], has played a central role in the development of low-dimensional topology over the last five decades [Man18, Hom23, Sav24]. The algebraic structure of $\Theta^{3}_{\mathbb{Z}}$ is linked to several topological problems such as the three fibers conjecture [Law88, Kol08], the disproof of the triangulation conjecture for topological manifolds in high dimensions [Man16], and so on. While it is known that the homology cobordism group contains an infinite rank summand by [DHST23], the existence of a torsion element is one of the most notable open problems about the structure of $\Theta^{3}_{\mathbb{Z}}$.
Seifert fibered spheres have been core objects to study the structure of $\Theta^{3}_{\mathbb{Z}}$. In particular, the subgroup $\Theta^{3}_{SF}$ generated by Seifert fibered spheres itself contains an infinite rank subgroup in $\Theta^{3}_{\mathbb{Z}}$ [Fur90, FS90]. Moreover, the aforementioned infinite rank summand [DHST23] lies also in $\Theta_{SF}^{3}$. On the other hand, there are several infinite families of Seifert fibered spheres known to bound contractible manifolds or homology balls [AK79, CH81, FS81, Fic84], namely trivial in $\Theta_{\mathbb{Z}}^{3}$. However, it is neither known in $\Theta_{\mathbb{Z}}^{3}$ whether there exists a non-trivial Seifert fibered sphere of finite order, nor, more generally, whether there exists any non-trivial linear dependence among them. More precisely, we ask:
@@
Let $\mathfrak{I}_{SF}$ be the image of $\Theta_{SF}^{3}$ under $h$. While $\mathfrak{I}_{SF}$ has been extensively studied in [DM19, DS19] and is known to be isomorphic to $\mathbb{Z}^{\infty}$, it is nevertheless possible to find many linear dependences in $\mathfrak{I}_{SF}$. Therefore, this fact and Question 1.1 motivate us to study the kernel of the restriction of $h$ to $\Theta^{3}_{SF}$. For example, the Brieskorn spheres $\Sigma(2,3,12n-7)$ for $n\geq 1$ are known to share the same local equivalence class [HM17], yet they were previously shown by Furuta [Fur90] to be linearly independent in $\Theta_{\mathbb{Z}}^{3}$.
@@
Therefore, potential counterexamples to Question 1.1 arise from linear dependence relations with respect to this explicit basis $\{h(B(n))\}_{n\geq 0}$ of $\mathfrak{I}_{SF}$. For any Seifert fibered sphere $Y$, there exists a unique homology sphere $Z_{Y}$ as the linear expression of $Y$ with respect to $\{B(n)\}$ such that $h(Y)=h(Z_{Y})$, or equivalently,
$Y\#-Z_{Y}\in\operatorname{ker}(h).$
We call such $Z_{Y}$ the associated homology sphere to $Y$ with respect to $\{B(n)\}$. For example, for any $Y$ in the aforementioned family $\Sigma(2,3,12n-7)$, the associated homology sphere $Z_{Y}$ is $\Sigma(2,3,5)=B(0)$.
In this paper, we find certain conditions for a Seifert fibered sphere $Y$ to guarantee that $Y$ and the associated homology sphere $Z_{Y}$ are not anymore related in $\Theta_{\mathbb{Z}}^{3}$. Such conditions are expressed in terms of the product $\mathcal{E}(Y)$ of the exponents of $Y$ and the Fintushel–Stern $R$-invariant [FS85]. Note that any linear combinations of the homology spheres of the form $Y\#-Z_{Y}$ are potential counterexamples to Question 1.1. We also prove their linear independence.
Theorem 1.2.
Let $Y$ be a Seifert fibered sphere and let $Z_{Y}$ be the associated homology sphere of the same image in $\mathfrak{I}$. If $R(Y)>0$ and $\mathcal{E}(Y)$ is distinct from $30$ and $(2m+1)(4m+1)(4m+3)$, then $Y\#-Z_{Y}$ is non-trivial in $\ker{h}$. Moreover, given such a family of homology spheres $Y(n)$ with the associated $Z(n)$, if all $\mathcal{E}(Y(n))$ are distinct, then Y(n)#Z(n)Y(n)\#-Z(n) are linearly independent in $\operatorname{ker}(h)$ up to rational homology cobordism.
Note that the previously known examples, such as $\{\Sigma(2,3,12n-7)\}$, also fall under Theorem 1.2. We remark that we obtain linear independence not only in $\Theta_{\mathbb{Z}}^{3}$, but also in the rational homology cobordism group $\Theta_{\mathbb{Q}}^{3}$ of rational homology spheres. Together with the assumption on $\mathcal{E}(Y)$ in Theorem 1.2, Furuta’s result [Fur90] is sufficient to prove linear independence in $\Theta^{3}_{\mathbb{Z}}$ of such homology spheres Y(n)#Z(n)Y(n)\#-Z(n). However, this does not directly imply linear independence in $\Theta_{\mathbb{Q}}^{3}$, since the kernel of $\Theta_{\mathbb{Z}}^{3}\rightarrow\Theta_{\mathbb{Q}}^{3}$ is known to be non-trivial [FS84, AL18, Sav20, Sim21].
Note that Theorem 1.2 indicates that the $d$-invariant [OS03a], the $\underline{d\mkern-2.0mu}\mkern 2.0mu$- and $\bar{d}$-invariants [HM17], and the $\phi_{n}$-invariants for all $n\geq 1$ [DHST23] must vanish for homology spheres Y(n)#Z(n)Y(n)\#-Z(n) in Theorem 1.2. In other words, for any homology sphere $Y$ obtained by a non-trivial linear combination of Y(n)#Z(n)Y(n)\#-Z(n), we see that:
$\underline{d}(Y)=d(Y)=\overline{d}(Y)=\phi_{n}(Y)=0.$
We observe that the $d$-invariant is not a complete rational homology cobordism invariant. One can find some homology spheres of infinite order in $\Theta^{3}_{\mathbb{Q}}$ with vanishing $d$-invariants from [NST24]. We also note that the $\bar{\mu}$-invariant [Neu80, Sie80] vanishes by the result of [DS19].
$\Theta^{3}_{\mathbb{Z}}\rightarrow\mathfrak{LE},$
where the latter group denotes the local equivalence group of the $\operatorname{Pin}(2)$-equivariant Seiberg–Witten Floer stable homotopy type. Recently, Dai, Sasahira, and Stoffregen [DSS23] proved that two Seifert fibered spheres have the same image in $\mathfrak{LE}$ if and only if they have the same image in $\mathfrak{I}$. Therefore, for a Seifert fibered sphere $Y$, the homology sphere $Y\#-Z_{Y}$ lies in $\operatorname{ker}(\Theta^{3}_{\mathbb{Z}}\rightarrow\mathfrak{LE})$ when the associated homology sphere $Z_{Y}$ to $Y$ is also Seifert fibered. Moreover, they also fall under Theorem 1.2 as well.
For those homology spheres, the $\kappa$-invariant [Man14] and the $\kappa o_{i}$-invariants for $i=0,\ldots,7$ [Lin15] from Seiberg–Witten Floer stable homotopy type, and the $\delta$-invariant [Fro10], the $\alpha$-, $\beta$-, and $\gamma$-invariants [Man16], and the $\underline{\delta}$- and $\overline{\delta}$-invariants [Sto17b] from Seiberg–Witten Floer homology must vanish. Namely, for any homology spheres $Y$ obtained by non-trivial linear combinations of Y(n)#Z(n)Y(n)\#-Z(n), we also have the following:
$\kappa(Y)=\kappa o_{i}(Y)=\alpha(Y)=\beta(Y)=\gamma(Y)=\underline{\delta}(Y)=\delta(Y)=\overline{\delta}(Y)=0.$
Based on [DS19], we construct explicit examples of Theorem 1.2 by taking $\{Y(n)\}$ as the families of the Brieskorn spheres:
@@
Their monotone graded subroots and hence their local equivalence classes were computed by Karakurt and the second author [KS22]. By using the fact that every monotone graded subroot is decomposed into simpler ones [DS19, Theorem 1.1], we first find associated homology spheres $Z_{1}(n)$ and $Z_{2}(n)$ to $Y_{1}(n)$ and $Y_{2}(n)$, respectively. From these linear relations, we obtain infinitely many homology spheres lying in $\operatorname{ker}(h)$ as follows:
$\big\{Y_{1}(n)\ \#\ -Z_{1}(n)\big\}_{n\geq 1}\cup\big\{Y_{2}(n)\ \#\ -Z_{2}(n)\big\}_{n\geq 1}.$
These families satisfy the conditions of Theorem 1.2 so that we obtain a substantial structural difference between $\Theta^{3}_{SF}$ and $\mathfrak{I}_{SF}$, providing affirmative evidence for Question 1.1. In other words, those families reprove an implicitly known result from other examples in several literature [Fur90, HM17, NST24].
Corollary 1.3.
Similarly, we provide examples of Theorem 1.2 whose local equivalence classes of Seiberg–Witten Floer stable homotopy type are also the same, by using the following family of the Brieskorn spheres:
@@
Relying on [KS22], we find their associated homology spheres $Z_{3}(n)$ that are Seifert fibered, so that the resulting homology spheres are as follows:
$\{Y_{3}(n)\ \#\ -Z_{3}(n)\}_{n\geq 1}.$
To the authors’ best knowledge, it is not known whether there exists a Seifert fibered sphere of the trivial local equivalence class but non-trivial in $\Theta^{3}_{\mathbb{Q}}$. The reason is that it is difficult to compute the $r_{s}$-invariant in general when the Fintushel–Stern $R$-invariant [FS85] is negative. Moreover, Issa and McCoy [IM18] proved that for a Seifert fibered sphere $Y$ with $d(Y)=0$, $R(Y)$ is always negative. Therefore, we conclude the introduction with the following question:
Question 1.4.
@@
Every manifold is assumed to be compact, connected, oriented, and smooth. The connected sum of $n$ copies of a manifold $Y$ is denoted by $nY$. We use the terms “homology sphere” and “homology ball” to refer to an “integral homology $3$-sphere” and an “integral homology $4$-ball”, respectively. $Y$ denotes a homology sphere unless otherwise stated. Seifert fibered spheres are oriented as the links of Brieskorn–Hamm complete intersection singularities in $\mathbb{C}^{3}$, so they bound plumbed $4$-manifolds with unique minimal negative definite plumbing graphs of central weights $e\leq-1$, see [Sav02, Section 1.1]. $\mathbb{F}$ denotes the field of characteristic $2$. Our grading shift convention for involutive Heegaard Floer invariants is in line with the one in [DS19].
@@ Acknowledgments
The authors would like to thank Marco Golla, JungHwan Park, and Masaki Taniguchi for helpful conversations. The authors are also grateful to Hayato Imori and Imogen Montague for reading the draft and providing detailed feedback. The first author is partially supported by the Samsung Science and Technology Foundation (SSTF-BA2102-02) and the NRF grant RS-2025-00542968. The second author was supported by the CNRS postdoctoral fellowship at the Laboratoire de Mathématiques Jean Leray in Nantes Université, France.
@@ 2. Preliminaries on Floer Theories
The proof of Theorem 1.2 involves two theories: involutive Heegaard Floer homology [HM17] and filtered instanton Floer homology [NST24]. In this section, we briefly recall the basic concepts, key properties, and essential notions of each theory. For the first two subsections, we review the involutive Heegaard Floer homology and its local equivalence class in terms of monotone graded subroots. In the third, we consider an analogue in the Seiberg–Witten Floer theory. In the last, we briefly recall the filtered instanton homology and the construction of the $r_{s}$-invariants.
@@ 2.1. Involutive Heegaard Floer homology and local equivalence group
$h:\Theta_{\mathbb{Z}}^{3}\rightarrow\mathfrak{I},\ \ h(Y)=(CF^{-}(Y)[-2],\iota),$
@@ 2.2. Monotone graded subroots for Seifert fibered spheres
In this subsection, we explain the roles of graded roots and their monotone graded subroots to describe local equivalence classes of Seifert fibered spheres. We follow the references [Nem05, DM19, DS19, KS22].
In [Nem05], Némethi introduced two combinatorial objects to compute Heegaard Floer homology of Seifert fibered spheres effectively. The first object is called a graded root $(R,\nu)$ equipped with a grading function $\nu$ on the vertex set of $R$, where $R$ is an upwards-opening tree with an infinite downwards stem defined from the negative definite plumbing graph for a Seifert fibered sphere. The second one is a graded $\mathbb{F}[U]$-module, called the lattice homology $\mathbb{H}^{-}(R)$. It is combinatorially defined from $(R,\nu)$, whose grading is induced by $\nu$ with the formal variable $U$ of grading 2-2. The grading on $\mathbb{H}^{-}(R)$ is denoted by the same symbol $\nu$ by abuse of notation. See [Nem05, DM19] for precise definitions.
@@
By combining the results of Ozsváth and Szabó [OS03b], Némethi [Nem05], and Dai and Manolescu [DM19], we have the following graded $\mathbb{F}[U]$-module isomorphism:
@@
where $\sigma$ denotes a grading shift [^1] For a Seifert fibered sphere, this shift is determined by algebraic topology of the minimal negative definite plumbing bounded by $Y$. See [KS22, Section 2.3] and references therein for a precise definition. and $\iota_{*}$ is the induced involution by $\iota$ on $HF^{-}(Y)$.
Following [DM19], we next construct a finitely generated free graded $\mathbb{F}[U]$-chain complex whose homology is the same as $\mathbb{H}^{-}(R)$. See [DM19, Lemma 4.3] for a proof of this fact.
@@
Then the generators of $C_{*}(R)$ are given as follows. For each leaf $v_{i}$, we place a single generator (also denoted by $v_{i}$ by abuse of notation) in grading $\nu(v_{i})$, so that we introduce an entire tower of generators $\mathbb{F}[U]v_{i}$. For each angle $\alpha_{i}$, we similarly place a single generator (denoted by also $\alpha_{i}$ by abuse of notation) in grading $\nu(\alpha_{i})+1$. We next define our differential to be identically zero on $v_{i}$, and set
@@
Let $(R,\nu,J)$ be a symmetric graded root, and let $\mathcal{G}_{R}$ be the image of the grading function $\nu:R\rightarrow\mathbb{Z}$. For a positive integer $n$, let $h_{1},\ldots,h_{n}$ (resp. $r_{1},\ldots,r_{n}$) be a sequence of decreasing (resp. increasing) even integers in $\mathcal{G}_{R}$ such that $h_{n}\geq r_{n}$. Then a monotone graded subroot
@@
By construction, the standard complex $(C_{*}(M),J)$ of a monotone graded subroot $M$ is a subcomplex of $(C_{*}(R),J)$, and it also forms an $\iota$-complex itself, see [DM19, Section 6]. However, this subcomplex is sufficient to describe the local equivalence class of a Seifert fibered sphere completely, as proven by Dai and Manolescu.
@@
Now we present a theorem, provided by Dai and Stoffregen [DS19], to describe how a monotone graded subroot decomposes into simpler ones up to local equivalence.
@@
For any Seifert fibered sphere $Y$, there exists a unique homology sphere $Z_{Y}$ expressed as the linear combination of the basis elements $\{B(n)\}$ such that $h(Y)=h(Z_{Y})$, i.e., the homology sphere $Y\#-Z_{Y}$ lies in $\operatorname{ker}(h)$. We call such $Z_{Y}$ the associated homology sphere to $Y$ with respect to $\{B(n)\}$. We omit the reference to the basis when it is clear from the context. In the final section of this article, beyond merely establishing the existence of $Z_{Y}$ and the linear independence of $Y$ and $Z_{Y}$ in $\Theta_{\mathbb{Q}}^{3}$, we will explicitly construct $Z_{Y}$ for a specific family $Y$ by applying Theorem 2.6.